NSF Proposal: RIGOROUS EXPERIMENTAL COMBINATORICS RIGOROUS EXPERIMENTAL COMBINATORICS: References Cited

[Aom] K. Aomoto, Jacobi polynomials associated with Selberg integrals, SIAM J. Math. Anal. 18(1987), 545-549

[AZ] A. Ayyer and D. Zeilberger, A bijectionist approach to the the Razumov-Stroganov conjecture, in preparation. (To be posted in arxiv.org and Zeilberger’s website)

[BB] J. Borwein and D. Bailey, Mathematics by Experiment. A.K. Peters, Natick, 2004. [BBCGLM], D. Bailey, J. Borwein, N. Calkin, R. Gingensohn, D. Luke, and V. Moll, ”Experimental Mathematics in Action”, A.K. Peters, 2007.

[BCG] E. Berlekamp, J. Conway, and R. Guy, “Winning Ways for your Mathematical Plays”, Academic Press, New York, 1982.

[Bour1] J. Bourgain, On triples in arithmetical progression, Geom. Funct. Anal. 9 (1999), 968-084.

[Bour2] J. Bourgain, Roth’s theorem on progression revisited J. Math. Anal. 104 (2008), 155-192.

[Bous] M. Bousquet-Mélou, “Rapport Scientifique pour obtenir l’Habilation à diriger des Recherches”,

avalable from http://dept-info.labri.u-bordeaux.fr/~bousquet/.

[Bre] D. M. Bressoud, “Proofs and Confirmations”, Math. Assoc. America and Cambridge University Press (1999).

[BS] E. Babson and E. Steingrímsson, Generalized permutation patterns and a classification of the Mahonian statistics, Sém. Loth. Comb. [http://mat.univie.ac.at/∼slc/] 44 , 2000, B44b, 18 pp.

[CS] Frédéric Chyzak and Bruno Salvy, Non-commutative elimination in Ore algebras proves multivariate identities, J. of Symbolic Computation 26 (1998), 187-227.

[DW] J. Désarménien and M. Wachs, Descent classes of permutations with a given number of fixed points, J. Comb. Ser. A 64 (1993), 311-328.

[Eri] J. Erickson, New Toads and Frogs Results, in: “Games of No Chance”, R. Nowakowski, ed., (Proc. of “Combinatorial Games at MRSRI, 1994”), 299-310, Cambridge University Press, 1996.

[Fis] I. Fischer, A new proof of the refined alternating sign matrix theorem, J. Combin. Theory Ser. A 114 (2007), 253-264.

[Foa1] D. Foata, On the Netto inversion number of a sequence, Proc. Amer. Math. Soc. 19 (1968), 236-240.

[Foa2] D. Foata, Distributions eulériennes et mahonniennes sur le groupe des permutations, in : “Higher Combinatorics”, M. Aigner, ed.,Berlin, 1976, 27-49.

[FT] M. Fisher and H. Temperley, Dimer Problems in Statistical Mechanics-an exact result, Philos. Mag. 6 (1961), 1061-1063.

[Ga] J. Gallian, Interview with Doron Zeilberger, MAA Focus May/June 2007, 14-17. (Available

on-line at www.maa.org).

[GG] A. Garsia and I. Gessel, Permutations Statistics and Partitions, Adv. in Math. 31 (1979), 288–305.

[GJ] I. Goulden and D.M. Jackson, An inversion theorem for cluster decompositions of sequences

C – 1

Doron Zeilberger

with distinguished subsequences, J. London Math. Soc.(2)20 (1979), 567-576.

[Gow] W.T. Gowers, A new proof of Szemerédi’s theorem, Geom. Funct. Anal. 11 (2001), 465-588.

[Kas] P. W. Kasteleyn, The statistics of dimers on a lattice: I. The number of dimer arrangements in a quadratic lattice, Physica 27 (1961), 1209-1225.

[KBI] V.E. Korepin, N.M. Bogoliubov and A.G. Izergin, “Quantum Inverse Scattering and Correlation Function”, Cambridge University Press, Cambridge, England, 1993.

[Ku1] G. Kuperberg, Another proof of the alternating sign matrix conjecture, Inter. Math. Res. Notes 1996, no. 3, 139-150.

[Ku2] G. Kuperberg, Symmetry classes of alternating sign matrices under one roof, Ann. of Math.

(2) 156 (2002), 835-866.

[Mac] P. A. MacMahon, Combinatory Analysis, vol. 1 and 2. Cambridge, Cambridge Univ. Press, 1915 (Reprinted by Chelsea, New York, 1955).

[NZ] J. Noonan and D. Zeilberger, The Goulden-Jackson Cluster Method: Extensions, Applications, and Implementations, J. Difference Eq. Appl. 5, 355-377, (1999).

[Oka] S. Okada, On the generating functions for certain classes of plane partitions, J. Comb. Theory, Series A 53 (1989), 1-23.

[Po1] G. Polya, “How to Solve it”, Doubleday, 1957.

[Po2] G. Polya, “Mathematics and plausible reasoning”, two volumes, Princeton Univ. Press, 1968.

[PWZ] M. Petkovsek, H. S. Wilf, and D. Zeilberger, “A=B”, AK Peters, Wellesley, (1996).

[RS] A.V. Razumov and Yu.G. Stroganov, Combinatorial nature of ground state vector O(1) loop model, Theor. Math. Physics 138 (2004), 395-400. math.CO/0104216.

[Sta1] R. Stanley Differentiably finite power series, European J. Combin. 1 (1980), 175-188.

[Sta2] R. Stanley, A baker’s dozen of conjectures concerning plane partitions. in: Combinatoire énumérative, G. Labelle and P. Leroux., eds., Lecture Notes in Mathematics 1234, 285-293. Springer Verlag, New York.

[Sta3] R. Stanley, Symmetries of plane partitions, J. Comb. Theory, Series A 43 (1986), 103-113.

[Ste] J. Stembridge, The enumeration of totally symmetric plane partitions, Advances in Mathematics 111, 227-243 (1995).

[SW] J. Sareshian and M. Wachs, q-Eulerian polynomials: excedance number and major index, Elec. Res. Announc. Amer. Math. Soc. 13 (2007), 33-45.

[SZ] B.Salvy and P. Zimmermann Gfun: a Maple package for the manipulation of generating and holonomic functions in one variable, ACM Trans. Math. Soft. 20(1994).

[Sze] E. Szemerédi, On sets of integers containing no k elements in arithmetic progressions, Acta Arith. 27 (1975), 299-345.

[Wei] E. W. Weisstein et al., “Ansatz.” From MathWorld–A Wolfram Web Resource.

http://mathworld.wolfram.com/Ansatz.html

[Wil] H. Wilf, Mathematics: An Experimental Science, in: “Princeton Companion of Mathematics”,

C – 2

NSF Proposal: RIGOROUS EXPERIMENTAL COMBINATORICS

T. Gowers, ed., 991-999, Princeton University Press, 2008.

[WZ] H. Wilf and D. Zeilberger, An algorithmic proof theory for hypergeometric (ordinary and “q”) multisum/integral identities, Invent. Math. 108, 575-633 (1992).

[Z1] D. Zeilberger, The Joy of Brute Force: The covariance of the number of inversions and the major Index, Personal Journal of Ekhad and and Zeilbeger,

http://www.math.rutgers.edu/~zeilberg/pj.html   .

[Z2] D. Zeilberger, A Holonomic systems approach to special functions identities, J. of Computational and Applied Math. 32, 321-368 (1990).

[Z3] D. Zeilberger, Proof of the alternating sign matrix conjecture, Elect. J. Combinatorics (http://www.combinatorics.org) 3(2) [Foata Festschrift] R13, (50 pages) (1996).

[Z4] D. Zeilberger, Proof of the refined alternating sign matrix conjecture, New York J. of Math. (elec.), (http://nyjm.albany.edu/) 2, 59-68 (1996).

C – 3